OpenStudy (anonymous):

find the slope of the line that passes though the pair of points (1,7),(10,1)

OpenStudy (anonymous):

\[\frac{y_2-y_1}{x_2-x_1}=\frac{1-7}{10-1}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

help

OpenStudy (anonymous):

it is arithmetic from there on in i am sure you can do the subtraction

OpenStudy (anonymous):

that would be 6/9 but thats not one of my answers

OpenStudy (anonymous):

no it is not for sure for example \(1-7\) is not \(6\)

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats still not one of my answers

OpenStudy (anonymous):

of course \[-\frac{6}{9}\] is not an answer choice either

OpenStudy (anonymous):

because that fraction is not in lowest terms you can reduce it

OpenStudy (anonymous):

to - 3/2?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i mean -2/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw